the square root of 5x+11=the sqaure root of 7x-1
Answer by venugopalramana(3286)

the square root of 5x+11=the sqaure root of 7x-1...SQUARING BOTH SIDES WE GET
5X+11=7X-1
7X-5X=11+1=12
2X=12
X=12/2=6